OpenStudy (anonymous):

please help me!! fan and medal!! 4. A wooded area in a state park has a mixture of different types of trees. There are 800 pine trees and 50 oak trees. The number of pine trees is decreasing at a rate of 5% per year. The number of oak trees is increasing at a rate of 15% per year. If these trends continue: A. Write two functions to model this situation, and graph those two functions on the same coordinate grid

OpenStudy (anonymous):

@sunshine201

OpenStudy (anonymous):

@uri @perl @samson245

OpenStudy (anonymous):

@sunshine201

OpenStudy (anonymous):

@PJtheVet

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

this is just like the compound interest formula the growth formula for Oak trees, A = future population, P = initial population, r is the growth rate n = number of years \[A = P(1 + \frac{r}{100})^n\] the pine trees are in decline... so its a depreciation formula \[A = P(1 - \frac{r}{100})^n\] the letters are identical to the growth situation... except the difference is growth contains a + and decline contains a - so substitute the information into each equation. so if you are having to graph them let y = A and x = n hope it helps

OpenStudy (anonymous):

So wait i have a question would i change the percents to decimals and put them for 4 and then plug in certain years?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@campbell_st i tried from years 1 to 10 and they never ever reach around the same amount of trees.

OpenStudy (campbell_st):

ok... so the simplified version of the 2 models are oak \[y = 50\times(1.15)^x\] pine \[y = 800 \times (0.95)^x\] so they are the 2 equations...in the 1st quadrant. graph the 2 equations... the point in intersection will be the time needed for the 2 tree populations to be equal in size